# Partner in Crime: Boosting Targeted Poisoning Attacks against Federated Learning

### v1

@@ IV-C Soft Label Design for Amplifier Set
Soft Label Design. Given the similarity between individual data samples, it is necessary to calculate the similarity between classes to design the appropriate soft labels for each class. With the data accessible to the attacker denoted as $\{S_{1},S_{2},\cdots,S_{N_{C}}\}$ and the converged surrogate model represented by $\boldsymbol{w}_{\textit{conv}}$, the class similarity between two classes, $c_{1}$ and $c_{2}$, based on feature distributions, can be represented as
@@
Instead of using binary values (0s and 1s), soft labels employ fractional values to capture the nuanced relationships between different data classes. From Equation (5), we can derive the class similarity between the source class and each intermediate class. Intuitively, to achieve the boosting objective, the portion of features in intermediate classes that are shared with the source class is maliciously labeled as the target, while the remaining features are correctly labeled. This approach is inspired by the label smoothing method [28], which involves using a uniform distribution over all classes to design soft labels for regularization. For example, consider an intermediate class with class ID $c_{z}$. The original hard label for $c_{z}$ is $\boldsymbol{e}_{c_{z}}=[0,0,\cdots,1,\cdots 0]^{T}$, where the $c_{z}$-th element is 1 and all the other elements are 0s. Here, $\boldsymbol{e}_{\text{src}}$ represents the source hard label, and $\boldsymbol{e}_{\text{tgt}}$ represents the target hard label. The soft label for $c_{z}$ is denoted as
$\boldsymbol{s}_{c_{z}}=CS_{\text{ftrs}}(c_{\text{src}},c_{z})*\boldsymbol{e}_{c_{\text{tgt}}}+(1-CS_{\text{ftrs}}(c_{\text{src}},c_{z}))*\boldsymbol{e}_{c_{z}}.$ (6)
If $CS_{\text{ftrs}}(c_{\text{src}},c_{z})$ is smaller than 0, $c_{z}$ retains its original hard label. In the final step of BoTPA, all data samples in the Amplifier set are modified to the crafted soft labels.
@@ IV-D2 Transition from vanilla attack to boosted attack
From Propositions 1 and 2, the vanilla and boosted poisoning effects steer the local model update in the directions of $\delta_{rs}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$ and $\delta_{rz}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$ , respectively, as shown in Equations (12) and (16). According to Equation (1), the intermediate class $z$ is selected based on its similarity to the source class $s$ in contributing to the model update. Additionally, in the actual computation, $-\log f_{\boldsymbol{w}}(\boldsymbol{x})$ is used to replace $f_{\boldsymbol{w}}(\boldsymbol{x})$ considering the cross-entropy loss function. Therefore, $\delta_{rz}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$ closely aligns with $\delta_{rs}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$, signifying that both the vanilla and boosted poisoning effects consistently drive the malicious models to update in analogous vector directions towards the adversarial objective, as shown in Figure 3. This indicates that BoTPA effectively enhances the impact of the targeted poisoning attacks while maintaining the underlying characteristics of the original attack vectors.
The boosting effect can accumulate when there are multiple training iterations locally within each communication round. Due to the enhanced poisoning effect in local models, the attacks can eventually influence the global model in favor of the adversarial objective. Overall, based on this theoretical analysis of weight divergence, we can attribute the boosting effect to the selection approach of intermediate classes. By accurately identifying and utilizing these intermediate classes, the attack can more effectively manipulate the model’s decision boundaries, thereby amplifying the impact of the poisoning efforts. This will be further demonstrated in Section VI-E.
@@ V Experimental Setup
V-A Datasets and FL systems
@@
Data Distributions. In IID scenarios (see Sections VI-A and VI-B), the dataset is uniformly and randomly assigned to each client. In non-IID scenarios (see Section VI-C), we adopt the setting from [48, 37, 16] to simulate data imbalance using a Dirichlet distribution [49], denoted as $\operatorname{Dir}_{K}(\beta)$. Here, $\beta$, ranging from 0 to positive infinity, represents the imbalance level, with smaller values indicating higher imbalance levels. In practice, $\beta$ is set to 0.5 and 1 to show BoTPA’s performance in scenarios with different levels of data imbalance. Figure 4 visualizes the non-IID data distributions using CIFAR-10.
Figure 4: Non-IID data distributions for CIFAR-10.
@@ V-C BoTPA Configurations
During the experimentation phase, the attacker constructs an Amplifier set using data samples from a fixed number of intermediate classes. Empirically, we set $N=2$ for FMNIST, $N=4$ for CIFAR-10, and $N=3$ for CH-MNIST, respectively. To determine the value of N, we incrementally increase its value from 1 in an “easy case” until we observe a decline in RI-ASR, choosing the value that yields the highest RI-ASR. The term “easy case” refers to the attack scenario involving a source-target combination that achieves a high RI-ASR in boosting data poisoning attacks (refer to Sections VI-A and VI-B for more details). This operation aims to balance targeted misclassification and minimize the impact on other classes. The performance of targeted poisoning attacks varies significantly across different source-target combinations and percentages of malicious clients. Consequently, it is not practical to identify a single optimal N across all class combinations due to the partial order nature of the optimization. Considering the slight variations in class-specific accuracy across different training rounds after convergence, we evaluate CIFAR-10 and FMNIST by averaging the ASRs from the 16th to the 25th rounds over three runs, and CH-MNIST by averaging the ASRs from the 30th to the 50th rounds.
@@ VI-A Targeted Data Poisoning Attack
TABLE II: Performance comparison of explicit poisoning attacks: ASRs (%) of vanilla and boosted attacks, along with the RI-ASRs (%).
    $\lambda=3$     $\lambda=5$
    V-ASR B-ASR RI     V-ASR B-ASR RI
FMNIST
Case 1     23.9 32.6 36.4     38.0 47.8 25.8
Case 2     32.5 39.6 21.8     50.4 61.9 22.8
CIFAR-10
Case 1     27.4 32.1 17.2     32.4 49.7 53.4
Case 2     24.0 28.0 16.7     39.3 40.7 3.6
CH-MNIST
Case 1     20.4 27.6 35.3     22.1 27.8 25.8
Case 2     13.6 18.2 33.8     15.3 22.4 46.4
Impact on Global Accuracy. The objective of our attack is to misclassify solely the source class while minimizing the impact on other classes, including intermediate ones. By monitoring global accuracy, we can assess the overall effect on other classes. Figure 5 presents the accuracy of global models for various datasets under vanilla and boosted attacks. For clarity, only scenarios with 10% and 30% malicious clients are presented. A detailed explanation of the selection strategy for the source-target combinations depicted in this figure will be provided in Section VI-B2. The results indicate a negligible difference in global accuracy between the vanilla and boosted attacks, implying that the boosted attack can take place in a stealthy manner without significantly affecting the accuracy of other classes.
@@ VI-B Targeted Model Poisoning Attack
VI-B1 Explicit Model Poisoning
We set the number of malicious clients to 10% and employ $\lambda=3$ and $\lambda=5$ to scale up the malicious updates, as the global model fails to ensure convergence for FMNIST and CIFAR-10 when $\lambda$ is increased further. Table II presents the ASRs for both vanilla and boosted attacks, as well as the RI-ASRs. For each dataset, the results are displayed for two source-target combinations, referred to as “case 1” and “case 2” for clarity. These cases are the median and best cases, respectively, derived from experiments with 10% malicious clients in data poisoning attacks. The results show that BoTPA can achieve RI-ASRs ranging from 3.6% to 46.4% when applied to explicit model poisoning attacks. This indicates that the boosted poisoning effect by BoTPA is further enhanced by the inherent model scaling in the vanilla attack framework. Similar to data poisoning attacks, the boosting effectiveness of BoTPA is not necessarily correlated with the inherent poisoning capability of vanilla attacks.
VI-B2 Stealthy Model Poisoning
@@
Similarly, Figures 7 and 8 compare the performance of vanilla and boosted attacks against the defenses of Median and Flame, respectively. The results of RI-ASRs show that the boosting performance is not strictly correlated with the percentage of malicious clients. Under these defenses, RI-ASRs are influenced by the trade-off between stealthiness and poisoning impact, which varies depending on the source-target pair. In most cases involving FMNIST and CIFAR-10, the highest RI-ASR is observed with 30% malicious clients, achieving up to 49.2% under Median and 33.3% under Flame. However, in some cases, such as CH-MNIST (class 0 → class 5), the boosting performance degrades as the number of malicious clients increases. This degradation occurs for two reasons: (1) the vanilla ASRs with fewer malicious clients are lower, making it easier to achieve higher RI-ASRs; (2) a larger proportion of malicious clients makes the model changes caused by mislabeling the Amplifier set more detectable and, consequently, more likely to be excluded. Overall, BoTPA demonstrates robust boosting performance across various datasets, despite performance variations among different source-target combinations.
@@ VI-E Feature Distributions in Latent Space
To understand how BoTPA affects the decision boundary between the source and target classes, we employ a visualization scheme based on penultimate layer representations, as proposed in [27]. This approach allows us to visually observe changes in the feature space. With the output from the penultimate layer, denoted as $P(\boldsymbol{x})$, the probability of an input $\boldsymbol{x}$ being classified into class $k$ is given by $f_{k}(x)=\frac{e^{P(\boldsymbol{x})^{T}\boldsymbol{w}_{k}}}{\sum_{i=1}^{N_{c}}e^{P(\boldsymbol{x})^{T}\boldsymbol{w}_{i}}}$, where ${\boldsymbol{w}_{i}}$ represents the weights connecting the penultimate layer and the last layer. The outputs of the penultimate layer are projected into 2D space in the following steps: (1) pick up 3 classes $\{c_{1},c_{2},c_{3}\}$ including the source and target class, and randomly sample 100 images for each class; (2) get the outputs of the penultimate layer for three classes $\{P(c_{1}),P(c_{2}),P(c_{3})\}$; (3) use the weights $\{\boldsymbol{w}_{1},\boldsymbol{w}_{2},\boldsymbol{w}_{3}\}$ to construct a set of orthonormal basis and project the penultimate layer outputs to the space; (4) apply Principal Component Analysis (PCA) to compress the space to 2D for visualization.
Figure 12 illustrates the latent feature space under vanilla and boosted poisoning attacks. For clarity, only the source and target classes are shown. We observe that (1) the source and target classes get closer in the latent feature space and (2) the clusters of source and target samples are more compact. This figure indicates that with BoTPA, data samples from the source class are more likely to cross the decision boundary and be misclassified as the target class.
@@ VII Potential Defense
VIII Conclusion & Future Work

### v2

@@ IV-C Soft Label Design for Amplifier Set
Soft Label Design. To design appropriate soft labels for each class, it is essential to calculate the similarity between classes. Let the data available to the attacker be denoted as $\{S_{1},S_{2},\cdots,S_{N_{C}}\}$, and the converged surrogate model as $\boldsymbol{w}_{\textit{conv}}$. The class similarity between two classes, $c_{1}$ and $c_{2}$, based on feature distributions, can be represented as
@@
Instead of using binary values (0s and 1s), soft labels employ fractional values to capture the nuanced relationships between different data classes. From Equation (6), we can derive the class similarity between the source class and each intermediate class. Intuitively, to achieve the boosting objective, the portion of features in intermediate classes that are shared with the source class is maliciously labeled as the target, while the remaining features are correctly labeled. This approach is inspired by the label smoothing method [34], which involves using a uniform distribution over all classes to design soft labels for regularization. For example, consider an intermediate class with class ID $c_{z}$. The original hard label for $c_{z}$ is $\boldsymbol{e}_{c_{z}}=[0,0,\cdots,1,\cdots 0]^{T}$, where the $c_{z}$-th element is 1 and all the other elements are 0s. Here, $\boldsymbol{e}_{\text{src}}$ represents the source hard label, and $\boldsymbol{e}_{\text{tgt}}$ represents the target hard label. The new label for $c_{z}$ is denoted as
$s_{c_{z}}=\begin{cases}e_{c_{z}},&\text{if }CS_{\text{ftrs}}(c_{\text{src}},c_{z})<0,\\ \text{Comb}(c_{src},c_{z}),&\text{otherwise},\end{cases}$ (7)
where $\text{Comb}(c_{\text{src}},c_{z})=CS_{\text{ftrs}}(c_{\text{src}},c_{z})\cdot e_{c_{\text{tgt}}}+\big{(}1-CS_{ftrs}(c_{\text{src}},c_{z})\big{)}\cdot e_{c_{z}}.$ This piecewise function is defined such that if $CS_{\text{ftrs}}(c_{\text{src}},c_{z})$ is smaller than 0, $c_{z}$ retains its original hard label. Otherwise, the label for $c_{z}$ is replaced with a soft label computed using the convex combination. In the final step of BoTPA, all data samples in the Amplifier set are modified to the crafted labels.
@@ IV-D2 Transition from vanilla attack to boosted attack
Analysis of Intermediate Class Selection: From Propositions 1 and 2, the vanilla and boosted poisoning effects steer the local model update in the directions of $\delta_{rs}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$ and $\delta_{rz}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$ , respectively, as shown in Equations (13) and (17). According to Equation (2), the intermediate class $z$ is selected based on its similarity to the source class $s$ in contributing to the model update. Additionally, in the actual computation, $-\log f_{\boldsymbol{w}}(\boldsymbol{x})$ is used to replace $f_{\boldsymbol{w}}(\boldsymbol{x})$ considering the cross-entropy loss function. Therefore, $\delta_{rz}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$ closely aligns with $\delta_{rs}(\boldsymbol{x},\boldsymbol{w}^{(k)}_{mT})$, signifying that both the vanilla and boosted poisoning effects consistently drive the malicious models to update in analogous vector directions towards the adversarial objective, as shown in Figure 3. This explains why BoTPA can effectively enhance the impact of targeted poisoning attacks.
Analysis of Soft Label Design: In Equations (16) and (17), the weight divergence is scaled by the factor $\lambda_{z}$, which originates from the soft label assigned to the intermediate class $z$. This indicates that, by utilizing soft labels, BoTPA strategically manipulates the feature representations of intermediate classes, positioning them as bridges between the source and target classes. This adjustment in the latent feature space reduces the separation of the decision boundary. Furthermore, compared to mislabeling intermediate classes via hard labels, the increased entropy in soft labels distributes the gradient contributions across the intermediate classes and the source class. This results in more subtle updates to the model parameters, thereby preserving the model’s normal behaviors on the intermediate classes.
The boosting effect can accumulate when there are multiple training iterations locally within each communication round. Due to the enhanced poisoning effect in local models, the attacks can eventually influence the global model in favor of the adversarial objective. Overall, the theoretical analysis of weight divergence underscores the critical role of accurately identifying and leveraging intermediate classes. This approach enables more effective manipulation of the model’s decision boundaries, thereby amplifying the impact of the poisoning efforts. This will be further demonstrated in Section VI-E.
@@ V Experimental Setup
V-A Datasets, FL systems and Adversarial Settings
@@
Data Distributions. In IID scenarios (see Sections VI-A and VI-B), the dataset is uniformly and randomly assigned to each client. In non-IID scenarios (see Section VI-C), we adopt the setting from [58, 43, 19] to simulate data imbalance using a Dirichlet distribution [59], denoted as $\operatorname{Dir}_{K}(\beta)$. Here, $\beta$, ranging from 0 to positive infinity, represents the imbalance level, with smaller values indicating higher imbalance levels. In practice, $\beta$ is set to 0.5 and 1 to show BoTPA’s performance in scenarios with different levels of data imbalance. [^2] Please refer to [43] for visualizations of the data distributions.
@@
Adversarial Settings: Poisoning attacks in FL can be broadly categorized into two types: data poisoning and model poisoning. Model poisoning attacks generally achieve higher success rates and greater stealthiness, while data poisoning attacks are less effective due to their limited attack capabilities. BoTPA, as an attack boosting method operating during the pre-training stage, is adaptable to various attack scenarios. To comprehensively evaluate its effectiveness, we apply it to both data poisoning and model poisoning scenarios. To determine the malicious client ratios, we follow the commonly adopted adversarial settings in existing attack and defense studies [17, 44]. Typically, the ratio of malicious clients starts at a low level. In this paper, we begin with 5%, corresponding to one or two clients in the FL systems across different datasets. This small ratio is chosen to evaluate the attack’s effectiveness under strict attack conditions. Then, we gradually increase the ratio to 30%, aligning with the Byzantine fault tolerance threshold of 33%, to assess the attack’s performance across varying attack scenarios.
@@ V-C BoTPA Configurations
During the experimentation phase, the attacker constructs an Amplifier set using data samples from a fixed number of intermediate classes. Empirically, we set $N=2$ for FMNIST, $N=4$ for CIFAR-10, and $N=3$ for CH-MNIST, respectively. To determine the value of N, we incrementally increase its value from 1 in an “easy case” until we observe a decline in RI-ASR, choosing the value that yields the highest RI-ASR. The term “easy case” refers to the attack scenario involving a source-target combination that achieves a high RI-ASR in boosting data poisoning attacks (see Sections VI-A and VI-B for more details). This operation aims to increase targeted misclassification and minimize the impact on other classes. The performance of targeted poisoning attacks varies significantly across different source-target combinations and percentages of malicious clients. Consequently, it is not practical to identify a single optimal N across all class combinations due to the partial order nature of the optimization. Considering the slight variations in class-specific accuracy across different training rounds after convergence, we evaluate CIFAR-10 and FMNIST by averaging the ASRs from the 16th to the 25th rounds over three runs, and CH-MNIST by averaging the ASRs from the 30th to the 50th rounds.
@@ VI-A Targeted Data Poisoning Attack
Impact on Global Accuracy. The objective of our attack is to misclassify solely the source class while minimizing the impact on other classes, including intermediate ones. By monitoring global accuracy, we can assess the overall effect on other classes. Figure 4 presents the accuracy of global models for various datasets under vanilla and boosted attacks. For clarity, only scenarios with 10% and 30% malicious clients are presented. A detailed explanation of the selection strategy for the source-target combinations depicted in this figure will be provided in Section VI-B. The results indicate a negligible difference in global accuracy between the vanilla and boosted attacks, implying that the boosted attack can take place in a stealthy manner without significantly affecting the accuracy of other classes.
@@ VI-B Targeted Model Poisoning Attack
Similarly, Figures 6 and 7 compare the performance of vanilla and boosted attacks against the defenses of Median and Flame, respectively. The results of RI-ASRs show that the boosting performance is not strictly correlated with the percentage of malicious clients. Under these defenses, RI-ASRs are influenced by the trade-off between stealthiness and poisoning impact, which varies depending on the source-target pair. In most cases involving FMNIST and CIFAR-10, the highest RI-ASR is observed with 30% malicious clients, achieving up to 49.2% under Median and 33.3% under Flame. However, in some cases, such as CH-MNIST (class 0 → class 5), the boosting performance degrades as the number of malicious clients increases. This degradation occurs for two reasons: (1) the vanilla ASRs with fewer malicious clients are lower, making it easier to achieve higher RI-ASRs; (2) a larger proportion of malicious clients makes the model changes caused by mislabeling the Amplifier set more detectable and, consequently, more likely to be excluded. Overall, BoTPA demonstrates robust boosting performance across various datasets, despite performance variations among different source-target combinations.
(a) FMNIST
(b) CIFAR-10
@@ VI-D Ablation Study
Here we conduct ablation experiments to evaluate the impact of different components in BoTPA. For this analysis, ASR values from vanilla attack scenarios are used as a baseline to clearly illustrate how different design choices affect the poisoning effect.
@@ VI-E Feature Distributions in Latent Space
To understand how BoTPA affects the decision boundary between the source and target classes, we employ a visualization scheme based on penultimate layer representations, as proposed in [33]. This approach allows us to visually observe changes in the feature space. Figure 11 illustrates the latent feature space under vanilla and boosted poisoning attacks. For clarity, only the source and target classes are shown. We observe that (1) the source and target classes get closer in the latent feature space and (2) the clusters of source and target samples are tighter. This figure indicates that with BoTPA, data samples from the source class are more likely to cross the decision boundary and be misclassified as the target class.
@@ VII Potential Defense
VIII Discussion
In this section, we analyze the computational complexity of BoTPA and discuss strategies to improve its scalability in FL systems with complex models or larger datasets. Finally, we explore its potential adaptation for real-world FL applications.
Computational Complexity: We assess computational complexity from three perspectives: training the surrogate model, selecting intermediate classes, and designing soft labels. For training the surrogate model, the complexity depends primarily on the size of the model and the size of the data set. In this study, we assume that the attacker utilizes all the data from compromised clients, and so the computational cost is determined by the malicious-client ratio and the number of data samples from these clients. Regarding the architecture of the surrogate model, it is important to note that the performance of the boosted attacks does not necessarily improve with deeper models, as shown in Figure 10. A simpler and shallower model that achieves performance comparable to the target FL model is preferable to reduce the training time overhead. For intermediate class selection and soft label design, the computational complexity in this step is affected by the number of intermediate classes, denoted as $\lvert C\rvert$. For a source class with $n_{s}$ samples and an intermediate class with $n_{i}$ samples, the complexity is proportional to $O(n_{s}\cdot n_{i})$. Considering all intermediate classes, the total complexity becomes $O(n_{s}\cdot\sum_{i=1}^{\lvert C\rvert}n_{i})$.
Scalability: To improve scalability in scenarios involving complex models and larger datasets, several strategies can be employed. First, instead of utilizing all samples from the source and intermediate classes, a random subset of samples can be selected to reduce computational overhead. From the perspective of model architecture, adopting a simpler surrogate model can significantly decrease the training time. Even if the surrogate model’s architecture differs slightly from the target model, it can still effectively capture meaningful class relationships. In addition, early stopping can be implemented to stop the training of the surrogate model once it achieves a predefined accuracy threshold, eliminating the need to wait for full convergence.
Attacking real-world FL systems: Designing effective attacks in FL systems for real-world domains, such as healthcare [61], presents several significant challenges. One critical issue is dealing with imbalanced datasets, which include class imbalance, biases in the data, and data sparsity. In real-world scenarios, implementing BoTPA requires adaptation to account for various types of data biases. When dealing with quantity and label distribution biases, the design of the Amplifier must consider differences in class size to ensure effective manipulation. For feature bias, individual data samples should be utilized in constructing the Amplifier, rather than grouping them by class, to better capture nuanced variations within the dataset.
IX Conclusion & Future Work
@@
Acknowledgments
We would like to thank the anonymous reviewers for their constructive comments, which help improve the quality of this article. This work was supported in part by the US Office of Naval Research under Grant N00014-23-1-2158 and the National Science Foundation under Grant CNS-2317829.